OpenStudy (raheelafzaal):

solve it with by parts x^2/(x+1)^1/2

OpenStudy (mathmale):

@raheelafzaal : I need more info. Are you attempting INTEGRATION? Do you mean, "integrate the following by parts"?

OpenStudy (mathmale):

"solving" this expression is not possible, because it's not an equation. However, you could certainly attempt "integration by parts" of this expression.

OpenStudy (raden):

u = x^2 ---> du = ... dv = 1/(x+1)^1/2 = (x+1)^-1/2 v = int (x+1)^-1/2 = ...

OpenStudy (raheelafzaal):

yes integrate